# Vinicius G. Goecks DEVCOM Army Research Laboratory Aberdeen Proving Ground, Maryland, USA

### 01 Feb 2024 v1

@@ V-F Results
Figure 7 presents a comparison of all baseline methods in terms of the average and standard deviation of the total rewards received during evaluation rollouts. COA-GPT, even in the absence of human interaction and relying solely on textual scenario representations, produces Courses of Action (COAs) that surpass the performance of all baselines on average. This includes the previously established state-of-the-art method, Autocurriculum Reinforcement Learning, which utilizes image data. COA-GPT-V, which also uses a single image as additional input, has equivalent performance compared to the previous baselines. Additionally, Figure 7 illustrates that the effectiveness of the COAs generated by COA-GPT is enhanced further (as indicated by a higher mean total reward) and exhibits reduced variability (evidenced by a lower standard deviation) when subjected to successive stages of human feedback. When taking into consideration the performance after human feedback, COA-GPT with vision models (COA-GPT-V+H1 and COA-GPT-V+H2) achieve higher mean total rewards compared to all previous baselines.
Figure 6 illustrates the evolution of generated COAs in response to human feedback. Initially, the COA generated without human input (Figure 6a) depicts the planned movement of friendly forces across bridges and their engagement with hostile units. After the first round of human feedback (“Make sure both our aviation units directly engages the enemy aviation unit.”), we see a strategic pivot (Figure 6b); the friendly aviation units are now directed to engage the enemy’s aviation assets directly. This adjustment reflects the human commander’s intent to prioritize air superiority. The second iteration of feedback — “Make sure only our Scout unit is commanded to control Bridge Bobcat (x: 75 y: 26) and our other assets (not the aviation) are split in two groups and commanded to move towards both enemy artillery using the attack_move command.” — as seen in Figure 6c), results in a more nuanced approach: the friendly forces are divided, with specific units tasked to target enemy artillery positions. Additionally, the reconnaissance unit is ordered to secure the northern bridge and the friendly aviation is still tasked to engage the threat aviation, demonstrating COA-GPT successfully followed the commander’s intent conveyed via textual feedback.
Figures 8 and 9 provide a comparative analysis of the mean and standard deviation of friendly and threat force casualties, respectively, during the evaluation rollouts. In Figure 8, we observe that the COA-GPT and COA-GPT-V, even when enhanced with human feedback (COA-GPT+H1, COA-GPT+H2, COA-GPT-V+H1 and COA-GPT-V+H2 models), exhibits higher friendly force casualties compared to other baselines. This outcome may be linked to COA-GPT’s lower control resolution, offering a single strategic command at the beginning of the episode rather than continuous command inputs throughout the episode as seen in the baseline methods. While this approach facilitates a more interpretable COA development process for human operators, it potentially increases casualty rates due to the limited tactical adjustments during COA execution.
@@ VI Conclusions
This research presents an advancement in the field of military planning and decision-making through the development of COA-GPT. In the face of increasingly complex and dynamic future battlefields, COA-GPT addresses a critical need in C2 operations for rapid and informed decision-making. By leveraging the power of LLMs, both with text-only and text and images as input, COA-GPT substantially accelerates the development and analysis of COAs, integrating military doctrine and domain expertise via in-context learning.
The results from comprehensive evaluations further validate the effectiveness of COA-GPT. Notably, it demonstrates superior performance in developing COAs aligned to commander’s intent, outperforming other methods including state-of-the-art Autocurriculum Reinforcement Learning algorithms. The enhanced adaptability and reduction in variability with human feedback highlight COA-GPT’s potential in real-world scenarios, where dynamic adaptation to changing conditions is critical.
Moreover, the ability of COA-GPT to generate actionable COAs within seconds, without the need for extensive pre-training, exemplifies its potential for rapid deployment in diverse military scenarios. This feature, coupled with its flexibility in adapting to new situations and the intuitive interaction with human commanders, underscores COA-GPT’s practical utility and operational efficiency.
In conclusion, COA-GPT represents a transformative approach in military C2 operations, facilitating faster, more agile decision-making and maintaining a strategic edge in modern warfare. Its development and successful application pave the way for further innovations in military AI, potentially reshaping how military operations are planned and executed in the future.

### 28 Mar 2024 v2

@@ V-B Quantitative Results
Figure 7 presents a comparison of all baseline methods in terms of the average and standard deviation of the total rewards received during evaluation rollouts. COA-GPT, even in the absence of human interaction and relying solely on textual scenario representations, produces Courses of Action (COAs) that surpass the performance of all AI baselines on average. This includes the previously established state-of-the-art method, Autocurriculum Reinforcement Learning, which utilizes image data. COA-GPT-V, which also uses a single image as additional input, has equivalent performance compared to the previous baselines. Additionally, Figure 7 illustrates that the effectiveness of the COAs generated by COA-GPT is enhanced further (as indicated by a higher mean total reward) and exhibits reduced variability (evidenced by a lower standard deviation) when subjected to successive stages of human feedback. When taking into consideration the performance after human feedback, COA-GPT with vision models (COA-GPT-V+H1 and COA-GPT-V+H2) achieve higher mean total rewards compared to all previous baselines. However, compared to the human expert, the best-proposed model achieves a 3.4% lower mean score, although it’s noteworthy that the human expert’s results demonstrate a larger variance.
Figures 8 and 9 provide a comparative analysis of the mean and standard deviation of friendly and threat force casualties, respectively, during the evaluation rollouts. In Figure 8, we observe that the COA-GPT and COA-GPT-V, even when enhanced with human feedback (COA-GPT+H1, COA-GPT+H2, COA-GPT-V+H1 and COA-GPT-V+H2 models), exhibits higher friendly force casualties compared to other baselines. This outcome may be linked to COA-GPT’s lower control resolution: offering a single strategic command at the beginning of the episode rather than continuous command inputs throughout the episode, as seen in all the baseline methods. While this approach facilitates a more interpretable COA development process for human operators, it potentially increases casualty rates due to the limited tactical adjustments during COA execution. The human expert, who was also able to continuously command the assets throughout the episode, achieved the lowest number of friendly casualties and the highest number of hostile casualties, on average.
@@ VI Conclusions
This research presents an advancement in the field of military planning and decision-making through the development of COA-GPT. In the face of increasingly complex and dynamic future battlefields, COA-GPT addresses a critical need in C2 operations for rapid and informed decision-making. Leveraging the power of LLMs, both with text-only and text and images as input, COA-GPT substantially accelerates the development and analysis of COAs, integrating military doctrine excerpts and domain expertise via in-context learning.
The results from the evaluations presented further validate the effectiveness of COA-GPT. Notably, it demonstrates superior performance in developing COAs aligned to the commander’s intent, closely matching expert humans and outperforming other AI methods, including state-of-the-art Autocurriculum Reinforcement Learning algorithms. While COA-GPT exhibits significant advantages in terms of speed and adaptability, it is enhanced further by the integration of human feedback. This collaboration leads to COAs that are aligned to human intent, yet are generated much more rapidly, underscoring the tool’s potential to transform the pace and efficiency of military planning.
Moreover, the synergy between COA-GPT and human commanders enriches the decision-making process, enabling the exploration of a broader range of operational possibilities and adjustments in real-time. This collaborative approach combines the speed and adaptability of AI with the nuanced understanding and strategic insight of human expertise, facilitating faster, more agile decision-making in modern warfare.
In conclusion, COA-GPT represents a novel approach to enhance military C2 operations. Its ability to generate actionable COAs within seconds, coupled with its flexible adaptation to new scenarios and intuitive interaction with human commanders, highlights its practical utility and potential for rapid deployment in diverse military scenarios. The development and successful application of COA-GPT pave the way for further innovations in military AI, potentially reshaping how military operations are planned and executed in the future, especially as it demonstrates an improved capacity for dynamic adaptation to changing conditions and enhanced performance through human-AI collaboration.
VII Future Work
Future work will extend the application of our methodology beyond the initial scenario presented, exploring its effectiveness across a diverse range of terrain and force configurations. The proposed method can be adapted to present commanders with multiple COA options, who can further iterate and refine these plans, or be adjusted to fit any established operational workflow. Additionally, future simulation platforms aim to develop a more complex model of opposition forces. This involves transitioning from the current simplistic representation, with fixed patrols and defensive positions, to dynamic simulations that capture the enemy commanders’ intentions and strategies, thereby enhancing the realism and strategic depth of our simulations.